# An Efficient Algorithm for Determining the Equivalence of Zero-one Reaction Networks

### v1

@@ 4 Algorithm
      Assume that at least two columns in $\mathcal{N}$ are the same. Without loss of generality, assume that the first $m_{1}$ columns of $\mathcal{N}$ are different and for any $i\in\{m_{1}+1,\dots,m\}$, there exists $j\in\{1,\dots,m_{1}\}$ such that $col_{i}(\mathcal{N})=col_{j}(\mathcal{N})$. By (21), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1}^{^{\prime}},\dots,a_{is}^{^{\prime}})^{\top}\in\mathbb{Q}^{s}$ such that
      $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}).$ (28)
      We denote by $\mathcal{N}_{1}$ and $\mathcal{N}^{^{\prime}}_{1}$ the first $m_{1}$ columns of $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$. Then, by (28), we have for any $i\in\{1,\dots,s\}$,
      $row_{i}(\mathcal{N}^{^{\prime}}_{1})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}_{1}).$
      Note that all the columns of $\mathcal{N}_{1}$ are different. Then, by (i), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$ such that
      $row_{i}(\mathcal{N}^{^{\prime}}_{1})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N}_{1}).$ (29)
      Note that for any $k\in\{m_{1}+1,\dots,m\}$, there exists $j\in\{1,\dots,m_{1}\}$ such that $col_{k}(\mathcal{N})=col_{j}(\mathcal{N})$. Hence, by (28), the $k$-th and the $j$-th coordinates of $row_{i}(\mathcal{N}^{^{\prime}})$ are equal. So, by (29), for any $i\in\{1,\dots,s\}$, $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N})$, where $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$.
  2. Case 2.
    Assume that some columns of $\mathcal{N}$ are not in the set $E$. By (21), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1}^{^{\prime}},\dots,a_{is}^{^{\prime}})^{\top}\in\mathbb{Q}^{s}$ such that
    $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\mathcal{N}).$ (30)
    Define two new matrices $\tilde{\mathcal{N}}$ and $\tilde{\mathcal{N}^{^{\prime}}}$ as follows. For any $j\in\{1,\ldots,m\}$, if $col_{j}(\mathcal{N})\in E$, then define
    $col_{j}(\tilde{\mathcal{N}}):=col_{j}(\mathcal{N}),\;\text{and}\;col_{j}(\tilde{\mathcal{N}^{^{\prime}}}):=col_{j}(\mathcal{N}^{^{\prime}}).$
    If $col_{j}(\mathcal{N})\not\in E$, then define
    $col_{j}(\tilde{\mathcal{N}}):=-col_{j}(\mathcal{N}),\;\text{and}\;col_{j}(\tilde{\mathcal{N}^{^{\prime}}}):=-col_{j}(\mathcal{N}^{^{\prime}}).$
    Then, by the definition of $E$, all the columns of $\tilde{\mathcal{N}}$ are in the set $E$. Note that by (30), for any $i\in\{1,\dots,s\}$,
    $row_{i}(\tilde{\mathcal{N}^{^{\prime}}})=\sum\limits_{j=1}^{s}a_{ij}^{^{\prime}}row_{j}(\tilde{\mathcal{N}}).$ (31)
    Then, by (31) and by (Case 1)–(ii), for any $i\in\{1,\dots,s\}$, there exists $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$ such that
    $row_{i}(\tilde{\mathcal{N}^{^{\prime}}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\tilde{\mathcal{N}}).$
    Therefore, by the definition of $\tilde{\mathcal{N}}$ and $\tilde{\mathcal{N}^{^{\prime}}}$, we have
    $row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{s}a_{ij}row_{j}(\mathcal{N}),$
    where $(a_{i1},\dots,a_{is})^{\top}\in S_{a}$.
Example 5.
We illustrate how Algorithm 3 works for $s=d=2$.
  1. Step 1
    Enumerate all possible zero-one reactions with two species (only enumerate one direction in a pair of reversible reactions).
    $0\xrightarrow{\kappa_{1}}X_{1},$ $0\xrightarrow{\kappa_{2}}X_{2},$ $0\xrightarrow{\kappa_{3}}X_{1}+X_{2},$
    $X_{1}\xrightarrow{\kappa_{4}}X_{2},$ $X_{1}\xrightarrow{\kappa_{5}}X_{1}+X_{2},$ $X_{2}\xrightarrow{\kappa_{6}}X_{1}+X_{2}.$
    Put the corresponding stoichiometric vectors (only consider different vectors) into a set:
    $E=\{(1,0)^{\top},(0,1)^{\top},(1,1)^{\top},(-1,1)^{\top}\}.$ (32)
  2. Step 2
    Let $S_{a}:=\emptyset$. Note that $d=2$. Choose $2$ column vectors from (32). There are six possibilities as follows:
    $\quad\quad\{\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\begin{pmatrix}1&-1\\ 0&1\end{pmatrix},\begin{pmatrix}0&1\\ 1&1\end{pmatrix},\begin{pmatrix}0&-1\\ 1&1\end{pmatrix},\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}\}.$
    For each matrix in the above set, do the following steps (we only show the details for the first matrix).
    1. Step 2.1.
      $\mathcal{A}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$.
    2. Step 2.2.
      By (22)–(24), we have
      $\Lambda_{1}=\Lambda_{2}=\emptyset,\Lambda_{3}=\{1,2\}.$
      Then, by (25), we have
      $F=$$\{(-1,-1)^{\top},(-1,0)^{\top},(-1,1)^{\top},(0,-1)^{\top},$
      $(0,0)^{\top},(0,1)^{\top},(1,-1)^{\top},(1,0)^{\top},(1,1)^{\top}\}.$
      For any $b\in F$, solve $a$ from $\mathcal{A}^{\top}a=b$. Then, we get the set of the solutions is
      $\{-1,0,1\}\times\{-1,0,1\}.$
      Add all these solutions into the set $S_{a}$.
    Similarly, we do Steps 2.1 and 2.2 for all other matrices. Then, we get the set $S_{a}$:
    $\{-1,0,1\}^{2}\cup\{(2,-1)^{\top},(2,1)^{\top},(-1,2)^{\top},(1,2)^{\top},$
    $(-1,-2)^{\top},(-2,-1)^{\top},(\frac{1}{2},\frac{1}{2})^{\top},(-\frac{1}{2},\frac{1}{2})^{\top},(\frac{1}{2},-\frac{1}{2})^{\top}\}.$
5 Implementation
We have implemented the procedure in Algorithm 1. The supporting codes are available online https://github.com/YueJ13/equivalence. As a larger example, we apply the algorithm to all the three-dimensional three-species five-reaction zero-one networks.
Step 1.
We enumerate all three-species five-reaction zero-one networks. There are $3819816$ networks.
Step 2.
We call CoefficientsForVacuous($3$, $3$), and by the computation carried out in Example 4, we get the set of rational vectors $S$ in (Step 2.). For each network from the $3819816$ networks, we do the following steps.
Step 2.1.
If the stoichiometirc matrix $\mathcal{N}$ of the network is rank three (i.e., the network is three-dimensional), and all rows of $\mathcal{N}$ are non-zero and change signs, then pick the network.
Step 2.2.
If the set $\mathcal{F}^{+}(\mathcal{N})$ is not empty, then pick the network.
Step 2.3.
If $c=3$ or $4$, then delete the network. (Note here, $c$ denotes the number of zero columns of the reactant matrix $\mathcal{Y}$.)
Step 2.4.
If there exists $a\in S$ such that
  • (1)
    $\mathcal{N}^{\top}a\neq\vec{0}$, and
  • (2)
    for any $i\in\{1,2,3\}$, if the $i$-th coordinate of $\mathcal{N}^{\top}a$ is non-zero, then $col_{i}(\mathcal{Y})=\vec{0}$,
then delete the network.
Then, we get $484477$ networks.
Step 3.
Delete the same networks after relabeling the species. Then, we get $76752$ networks.
Step 4.
We call CoefficientsForEquivalence($3$, $3$), and we get a set of vectors $S_{a}\subset{\mathbb{Q}}^{3}$, where for any $a=(a_{1},a_{2},a_{3})\in S_{a}$, we have
$a_{i}\in\{\pm 4,\pm 3,\pm 2,\pm\frac{3}{2},\pm\frac{4}{3},\pm 1,\pm\frac{3}{4},\pm\frac{2}{3},\pm\frac{1}{2},\pm\frac{1}{3},\pm\frac{1}{4},0\}.$
For each pair of networks $G$ and $G^{^{\prime}}$ from the $76752$ networks in Step 3, check if $\mathcal{N}$ and $\mathcal{N}^{^{\prime}}$ (after permutation according to the description in Step 8.1 of Algorithm 1) satisfy the following condition: for any $i\in\{1,2,3\}$, there exist $(a_{1},a_{2},a_{3})^{\top}\in S_{a}$ and $(b_{1},b_{2},b_{3})^{\top}\in S_{a}$ such that
$\qquad\quad\enspace row_{i}(\mathcal{N}^{^{\prime}})=\sum\limits_{j=1}^{3}a_{j}row_{j}(\mathcal{N}),\;row_{i}(\mathcal{N})=\sum\limits_{j=1}^{3}b_{j}row_{j}(\mathcal{N}^{^{\prime}}).$
If yes, then put the networks $G$ and $G^{^{\prime}}$ into the same class. Then, we get $40522$ classes of networks. Choosing the first network from every class, we get $40522$ networks.
@@ 5 Implementation
Calculate the reduced Gröbner bases of the $40522$ networks. Delete the networks whose reduced Gröbner bases are $1$ or have a monomial as an element. Then, we get $32668$ networks.
@@
Combine the networks having the same reduced Gröbner basis into the same class. Choosing the first network from every class, we get $32394$ networks.
We record the timings for carrying out the above steps in the following Table 1. In the experiments, we use the command Basis [13] in Maple [22] to calculate the reduced Gröbner bases. Note that if we do not apply Algorithm 1, then we need about $60$ hours to check the reduced Gröbner bases of all three-dimensional three-species five-reaction zero-one networks. Our purpose is to show the improvement Algorithm 1 makes comparing with purely computing Gröbner bases. One can improve the computational timings presented in Table 1 by applying any more advanced tools for computing Gröbner bases rather than the standard Maple package.
Table 1: Computational time (min: minutes; s: seconds)
@@
Step 1 3819816 1 min
Step 2 484477 30 min
Step 3 76752 145 s
Step 4 40522 40 min
Step 5 32668 100 s
Step 6 32394 30 min
@@
Acknowledgments We thank Professor Balázs Boros for his suggestion on how to efficiently generate small networks in the Joint Annual Meeting KSMB-SMB 2024. We thank Professor Murad Banaji for generating and providing all three-dimensional three-species five-reaction zero-one networks by his own efficient procedure as a reference for us.

### v2

@@ 5 Implementation
    for any $i\in\{1,2,3\}$, the $i$-th coordinate of $\mathcal{N}^{\top}a$ is non-zero implies $col_{i}(\mathcal{Y})=\vec{0}$,
then delete the pair.
Then, we get $91802$ networks.
@@
Combine the remaining pairs $\{\mathcal{N},\mathcal{Y}\}$ into the same group if they have the same reduced row echelon form of $\mathcal{N}$. For each group, we do the following steps.
Step 5.1.
For each pair $\{\mathcal{N},\mathcal{Y}\}$ and $\{\hat{\mathcal{N}},\hat{\mathcal{Y}}\}$, we do the following steps.
Step 5.1.1.
Check if there exists a permutation on the rows of $\mathcal{Y}$ such that $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ are equal after permuting the rows of $\mathcal{Y}$.
Step 5.1.1.1.
If yes, then put the pair of matrices $\{\mathcal{N},\mathcal{Y}\}$ and $\{\hat{\mathcal{N}},\hat{\mathcal{Y}}\}$ into the same group.
Step 5.1.1.2.
If not, then consider the next pair of matrices.
Step 5.2.
For every group, choose the first pair $\{\mathcal{N},\mathcal{Y}\}$ as the representative element.
Then, we get $48703$ networks (pairs).
@@
For every remaining network, calculate the reduced Gröbner bases for steady-state ideal $\langle f\rangle$. Delete the networks whose corresponding reduced Gröbner bases are $\{1\}$ or have a monomial as an element. Then, we get $38249$ networks.
Step 7.
Combine the networks with the same reduced Gröbner basis into the same class. Choose the first network as the representative element. Finally, we get $38185$ networks.
We record the computational results for the above steps in Table 1. In the experiments, we use the command Basis [14] in Maple [20] to calculate the reduced Gröbner bases. Note that the number of all three-dimensional three-species five-reaction zero-one networks is $3819816$ [16]. If we do not apply Algorithm 1, then we need about $60$ hours to check the reduced Gröbner bases of these networks. Our main purpose is to demonstrate the improvements achieved by Algorithm 1 compared with the pure computation of Gröbner bases. The computational time of Step 7 presented in Table 1 can be improved by applying any more advanced tools to calculate the Gröbner bases rather than the standard Maple package. Recall that the original version of Algorithm 1, as presented in [16], requires approximately one hour to complete the entire process (see [16, Table 1]). Here, the improved version has saved about half of the total computational time by operating on stoichiometric matrices rather than networks in the first three steps. Another observation is that most inconsistent networks are filtered out when the vacuous ideals are verified at Step 4.
Table 1: Computational results for $(3,5,3)$-networks (min: minutes; s: seconds)
STEP NUMBER OF Matrices TIME
Step 1 43520 1 min
Step 2 14444 1 min
Step 3 2548 1 min
@@
Step 4 91802 1 min
Step 5 48703 30 s
Step 6 38249 2 min
Step 7 38185 22 min
@@
In addition, we apply Algorithm 1 to two key classes of zero-one networks: three-dimensional (with $3$ species and $6$ reactions) and four-dimensional (with $4$ species and $5$ reactions). Given that both classes include a large number of networks (the numbers of matrices at Step 1 are $276382$ and $7818144$ respectively), we restrict our computations to quadratic networks only. Notably, Algorithm 1 can be easily adjusted to compute quadratic zero-one networks (in Step 1, we only enumerate the stoichiometric vectors for the quadratic reactions and in Step 4, we only consider the reactant matrices $\mathcal{Y}$} for quadratic networks). We record the timings for the completion of the steps of Algorithm 1 in Table 2 and Table 3, respectively. Another interesting finding is that there exists no networks after carrying out Step 6 for these two classes of quadratic networks. That means these networks have no positive steady states for a generic choice of rate constants. We show an example of three species six-reaction three-dimensional zero-one quadratic network remaining in Step 5, see Example 7.
Table 2: Computational results for $(3,6,3)$-quadratic networks (min: minutes; s: seconds)
STEP NUMBER OF Matrices TIME
Step 1 10693 1 min
Step 2 5628 1 min
Step 3 960 1 min
STEP NUMBER OF NETWORKS TIME
Step 4 1897 1 min
Step 5 1565 0.2 s
Step 6 0 0.14 s
Step 7 NULL NULL
Table 3: Computational results for $(4,5,4)$-quadratic networks (min: minutes; s: seconds)
STEP NUMBER OF Matrices TIME
Step 1 1440168 3 min
Step 2 112008 1 min
Step 3 4822 5 min
STEP NUMBER OF NETWORKS TIME
Step 4 21440 1 min
Step 5 5032 1.9 s
Step 6 0 2.9 s
Step 7 NULL NULL
Example 7.
Consider the following $3$-dimensional quadratic zero-one network with $3$ species and $6$ reactions:
$X_{2}+X_{3}\xrightarrow{\kappa_{1}}X_{1}+X_{3},$ $X_{1}+X_{3}\xrightarrow{\kappa_{2}}X_{2}+X_{3},$ $X_{1}+X_{3}\xrightarrow{\kappa_{3}}X_{1}+X_{2},$
$X_{1}+X_{2}\xrightarrow{\kappa_{4}}X_{1},$ $X_{1}+X_{2}\xrightarrow{\kappa_{5}}X_{2}+X_{3},$ $X_{1}+X_{2}\xrightarrow{\kappa_{6}}X_{1}+X_{2}+X_{3}.$
The steady-state system $f$ is
$f_{1}$ $=\kappa_{1}x_{2}x_{3}-\kappa_{2}x_{1}x_{3}-\kappa_{5}x_{1}x_{2},$
$f_{2}$ $=-\kappa_{1}x_{2}x_{3}+\kappa_{2}x_{1}x_{3}+\kappa_{3}x_{1}x_{3}-\kappa_{4}x_{1}x_{2},$
$f_{3}$ $=-\kappa_{3}x_{1}x_{3}+\kappa_{5}x_{1}x_{2}+\kappa_{6}x_{1}x_{2}.$
From $f_{1}=f_{2}=f_{3}=0$, we can get
$f_{1}+f_{2}+f_{3}=(-\kappa_{4}+\kappa_{6})x_{1}x_{2}=0.$
Hence, if $\kappa_{4}=\kappa_{6}$, the network admits infinitely many positive steady states that satisfy
$x_{1}=\frac{\kappa_{1}\kappa_{3}}{\kappa_{3}\kappa_{5}+\kappa_{2}\kappa_{5}+\kappa_{2}\kappa_{6}}x_{3},\;x_{2}=\frac{\kappa_{3}}{\kappa_{5}+\kappa_{6}}x_{3}.$
However, the reduced Gröbner basis of the steady-state ideal $\langle f\rangle$ in ${\mathbb{Q}}(\kappa)[x]$ is
$\{x_{2}x_{3},\ x_{1}x_{3},\ x_{1}x_{2}\}.$
That means the network indeed has no positive steady states for any generic choice of $\kappa\in{\mathbb{R}}_{>0}^{6}$. So, this network is consistent and has a non-vacuous steady-state ideal—hence its passage through Steps 2 and 4 of Algorithm 1—yet it is still dynamically trivial.
Acknowledgements. We thank Professor Balázs Boros for his suggestion on how to efficiently generate small networks in the Joint Annual Meeting KSMB-SMB 2024. We thank Professor Murad Banaji for generating and providing all three-dimensional three-species five-reaction zero-one networks by his own efficient procedure as a reference for us. We would like to thank the anonymous reviewers of ISSAC 25 for their insightful and valuable comments, which have greatly contributed to the improvement of this work.